# An Industry Interview Study of Software Signing for Supply Chain Security

### v1

@@ 4.3 Data Collection
Our Institutional Review Board (IRB) approved this study.
@@
TABLE II: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives. “Technical leader” — senior, lead, partner, and principal engineers. “Engineer” and “Manager” — junior staff.
ID Role Experience Product Area
S1 Research leader 5 years Digital technology
S2 Senior mgmt. 15 years SSC security
S3 Senior mgmt. 13 years SSC security
S4 Technical leader 20 years Cloud, Dev tools
S5 Engineer 2 years Cloud security
S6 Technical leader 27 years Digital technology
S7 Manager 6 years Cloud security
S8 Technical leader 8 years Internet services
S9 Engineer 2.5 years SSC security
S10 Engineer 13 years Dev tools
S11 Technical leader 16 years Social technology
S12 Technical leader 4 years Cloud, OSS security
S13 Senior mgmt. 16 years Telecommunications
S14 Research leader 13 years Cloud security
S15 Senior mgmt. 15 years Aerospace security
S16 Technical leader 26 years Digital technology
S17 Senior mgmt. 15 years SSC security
S18 Manager 11 years Internet services
@@
Target Population: To address our research questions effectively, the target population was industry practitioners with expertise in software signing. An ideal subject would have in-depth knowledge and experience related to software supply chain practices within their organizations. Such experts would have high credibility, insider perspectives, and nuanced interpretations.
Recruitment: We employed a non-probability-based Purposive and Snowball population sampling approach to recruit such subjects [82]. Our initial subject sample pool was members from the Cloud Native Computing Foundation’s Kubecon 2023 conference organizing committee, many of whom also contribute to software supply chain security projects hosted by The Linux Foundation. This pool contained many security experts, although we acknowledge that the connection to specific signing projects may bias their responses to this genre of signing approach. Subsequently, we expanded our subject pool through recommendations from this initial group of subjects and other industry contacts. We offered each participant a \$100 gift card as an incentive.
@@
In total, we contacted 30 candidates, of whom 18 agreed to participate (60% response rate). These subjects came from 13 different organizations. Their demographic information is summarized in Table II. Subjects were affiliated with companies with security products or were part of their company’s security team. For subject anonymity, we do not distinguish those from CNCF or from snowballing.
@@ 4.4 Data Analysis
Interview recordings were transcribed by www.rev.com (human transcription). We anonymized identifying information. Our analysis then proceeded in several stages.
(Stage 1) We began with data familiarization [85]. Three analysts (A1–A3) undertook an inductive reading, memoing, and coding process for all transcripts.
(Stage 2) Subsequently, following O’Connor & Joffe’s recommendation [86] to initially independently code 10-25% of the data, we randomly selected approximately 33% (6 transcripts) for multiple memoing. Two analysts (A1, A2) conducted this multiple memoing of data in three rounds, collaborating to discuss emerging code categorizations after each round. This process yielded 506 coded memos, which were reviewed by Analysts A1 and A3 over four one-hour meetings to create a codebook comprising 36 code categories, excluding demographics.
(Stage 3) To ensure the reliability of our coding, analysts A1 and A3 randomly selected and coded a transcript. We used the Percentage agreement metric as recommended by Feng [87] for easy coding tasks with nominal data. Since our codebook was collectively defined by both analysts A1 and A3, we assumed this coding task should not be too difficult — we achieved an 89% agreement rate, confirming this assumption. Any discrepancies in coding were resolved through discussion and code redefinition.
(Stage 4) Next, a second set of 6 transcripts was selected and coded by Analyst A1 using an initial codebook, adding 6 minor categories, each with less than 6 codes. This updated codebook was reviewed by Analyst A3.
@@
(Stage 5) The same process was applied to the final 6 transcripts, yielding 4 more minor categories with fewer than 4 codes each. Out of the 10 new categories, only 2 provided new insights into our data, as determined through multiple review rounds by Analysts A1 and A3.
@@
(Thematic and Framework Analyses) From this process we induced themes following Braune & Clark’s method [88]. These themes helped us answer RQs 2-4 directly. To answer RQ1, we also applied a framework analysis [89], mapping the themes to the software supply chain factory model as a reference framework to analyze participants’ responses and further categorize the themes.
@@
Through our coding phase’s thematic analysis, we identified eight themes. Four of these themes were focused on the practical implementation of Signing, while the remaining four addressed various aspects of software signing: the challenges encountered, its perceived importance, the impact of security failures on its adoption, and the influence of security standards and regulations on its adoption.
@@
To determine whether additional aspects of the phenomenon might be expected to emerge if we continued to sample, we measured saturation in the 18 interview transcripts. Following Guest et al.’s [90] recommendation, we tracked the cumulative appearance of new codes in each interview. Each interview had a median of 33 unique codes, also depicted in  Figure 5. Saturation was achieved after interview 11 (details in Appendix Appendix D).
4.5 Study Limitations
@@
Positionality Statement: We acknowledge the role that our backgrounds may have played in this study [91]. The author team are cybersecurity researchers with expertise in software signing, so our expertise both enabled and influenced the protocol development and analysis. Beyond this, one of the authors is a contributor to a popular signing tool. To reduce bias in our study, that author had no involvement in the development of the research protocol or analysis. However, that author reviewed the results and framing of the study, to increase its validity for the software signing community.
Other Threats to Validity: This work is qualitative, and bears the limitations of any qualitative study. First, the cost of data collection reduces the sample size, threatening generalizability. Our subject recruitment may also have introduced biases, e.g., towards CNCF/The Linux Foundation-oriented frameworks and tools. As mitigations, we note that our codebook saturated at the 11th interview, with few/no new codes observed in subjects 12-18 (Appendix D); and our sample size is typical for similar studies [71, 72]. Second, qualitative data implies subjective analysis. To improve the reliability of our results, we followed an iterative process to develop the codebook and used multiple raters with a high measure of agreement.
Specific to the topic of software signing, the frameworks for software supply chain security and the available automation for implementing signing are evolving. Our results should be understood as a snapshot of current practices in industry. They will grow out of date as signing standards, regulations, and tooling continues to evolve.
@@ 5.1 RQ1: Signing Implementations in Practice
Finding 1. We refine the software supply chain factory model from Figure 1, as shown in Figure 4 and Table III. Our refined model shows candidate points for the establishment of provenance. We analyze our subjects’ responses using this refined model, identifying points on this model where subjects’ teams or organizations employ software signing.
Figure 4: Our refined software supply chain factory model (compare to Figure 1), showing different points where provenance and integrity of software could be established (PI, PE, PB, PD, PS, PA) and Verified (VI, VE, VB, VD, VS, VA) using software signatures. Abbreviations are defined in Table III.
TABLE III: Different Signing use Points as noted in  Figure 4 and subjects who employ software signing at these points. We mark subjects who consider them optional for reasons of demand by customers, nature of products, etc.
Point Description Subjects
@@
PI Internal code contributors sign their commits S2-S11, S14-S17 (S4, S8, S14 – Optional)
PE External code contributors sign their commits S7, S10 (S11, S14 — Optional)
VI Verification of signatures from Internal contributors S7, S14, S17
VE Verification of signatures from external contributors
PSC Signing after code reviews/audits S2-S11, S14-S17 (S4, S8, S14 — Optional)
@@ VI Verification of signatures from Internal contributors S7, S14, S17
VSC Verification of source code signatures before build S2, S3, S7, S14
PB Signing of build output S1-S5, S7, S8, S11-S15, S17 (S8 —Optional)
@@
VB Verification of build signature before Deployment or final artifact S7, S14
PA Signing of Final Software product All S1–S18
@@
PD Signing after verification and certification of dependencies S6, S7, S9, S16, S17, S18
VD Verification of Signatures from external dependencies S9, S11, S17 (S9, S11, S17 — Optional)
VA Verification by final customers Difficult to determine
Using our initial analysis plus the recommendations and philosophy of several frameworks, which recommend a continuous verification of integrity and trust in code alone (not humans) [20, 92], we present a refined form of the software supply chain factory model as shown in Figure 4. Our model comprehensively identifies points in the software delivery process where the security property of provenance is expected to be established or verified. We compare this with subjects’ reported usage of other security methods for these factory model components, and second with framework (and standards) prescriptions.
@@ 5.1.1 Source code
First, we ascertained whether our subjects used any form of external code contributors in their software creation process. Out of 10 subjects who reported a form of collaboration with open-source or other external contributors, only 5 subjects used external code contributions in the production code of projects they owned (both commercial and OSS). We then proceeded to determine if any form of signing is used to establish provenance for points PE, PI, VE, VI, and PSC.
External contributors sign commits (PE): Signing is mostly required by teams who use external code contributors. Out of 5 subjects whose teams used external code contributions in their production code, 4 of these subjects require that commits be signed. However, two of those do not require signing for open-source projects or noncritical projects. Only 2 of our subjects unconditionally require signing from their external contributors. As S2 put it, “I got something signed from somebody I don’t know…So we do the code review, and then we have a sign-off by somebody on our team that signs off from that code”.
Internal contributors sign commits (PI): Signing is almost always required by teams for their internal code contributors. 14 of 18 subjects report that their team required some sort of commit signing from internal contributors. Subjects who list this as an optional requirement (3) cite similar reasons as in PE above.
Verification of signatures from contributors (VE&VI): While signing is relatively required, it is not verified as a check for security and provenance of code contributions from contributors. Techniques like code reviews (and security audits), security/vulnerability scans, tests, etc, are the most used security methods for assessing the security of code contributions. S10 opined to this, “I think everybody on the team does sign their commits, but we don’t really verify it…If somebody stopped signing their commits, I don’t think much would really happen”
Code is signed after reviews/audit (PSC): Signing here is common similar to PI. Subjects who report that signing is required in PI, also report that all activities related to code reviews, or audits are (cryptographically) signed off on.

### v2

@@ 4.3 Data Collection
Table 2: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff. We also highlight type of software subject’s immediate team produces. * refers to cases where subject’s team built other types of software in addition to the main stated. O refers to cases where subject belonged to multiple teams with other products to the one stated.
ID Role Experience Software Type
S1 Research leader 5 years Internal POC software
S2 Senior mgmt. 15 years SAAS security tool*O
S3 Senior mgmt. 13 years SAAS security tool*O
S4 Technical leader 20 years Open-source tooling
S5 Engineer 2 years Internal security tooling
S6 Technical leader 27 years Internal security tooling*O
S7 Manager 6 years Security tooling*O
S8 Technical leader 8 years Internal security tooling*
S9 Engineer 2.5 years SAAS security*
S10 Engineer 13 years SAAS security*
S11 Technical leader 16 years Firmware*
S12 Technical leader 4 years Consultancy
S13 Senior mgmt. 16 years Internal security tool
S14 Research leader 13 years POC security Software*
S15 Senior mgmt. 15 years Internal security tooling
S16 Technical leader 26 years Security tooling
S17 Senior mgmt. 15 years SAAS
S18 Manager 11 years Security tooling
@@
Target Population: To effectively address our research questions, the target population was industry professionals with expertise in software signing. An ideal subject would have deep knowledge and experience related to software supply chain practices within their organizations. Such subjects are typically middle- or high-ranking engineering staff.
Recruitment: Given the seniority and expected workload of the target population, we anticipated recruitment to be a challenge. To mitigate this, we employed a non-probability-based purposive and snowball sampling approach to recruit participants [10]. This approach incurs the risk of bias, but has the reward of access. Given the understudied topic at hand, we felt that the reward outweighed the risk.
Our initial sample pool of subjects was made up of members of the Kubecon 2023 conference organizing committee of the Cloud Native Computing Foundation. This group matches our target population: all of them either worked as a member of the software security teams of their organizations or worked in organizations with products and services in the domain of software security, and all were medium- to high-ranking members of their organizations. Many of these individuals also contribute to software supply chain security projects hosted by The Linux Foundation, making them well-suited to provide expert insights. We had special access to this group because one of the authors was a member of that committee. However, recognizing the potential biases of this initial sample, we expanded our recruitment through recommendations from initial participants and other industry contacts. This approach allowed us to diversify our subject pool and gather a broader range of insights, while still focusing on individuals with substantial expertise in the domain.
@@
In total, we contacted 30 candidates, of whom 18 agreed to participate (60% response rate). These subjects came from 13 different organizations. The subjects were affiliated with companies with security products or were part of their company’s security team. Each participant was offered a \$100 gift card as an incentive, in recognition of their high rank.
Subject Demographics: We summarize the demographic and organization information of the subjects in Table 2 and  Table 6. We note that six participants were from the initial KubeCon committee invitations; industry contacts of the authors yielded seven participants; and snowball sampling (recommendations from participants) yielded five participants.
In addition to the highlighted, we also note that 13 subjects (from 11 organizations – 4 Small, 2 Medium, and 6 Large organizations) confirmed that they either lead, or belong to, teams involved in software supply chain, infrastructure and tooling, or security control implementation. These subjects indicated that they are responsible for initiating or implementing their organization’s security controls and have responsibilities related to compliance. Although not all subjects directly owned compliance, their involvement in security strategy and operational decision making provided them with the appropriate context to assess the utility, influence behind adoption, and challenges affecting the implementation of Software signing.
Subjects from small organizations typically focused on SAAS products related to software supply chain, container, and cloud security; or were consultants. In medium-scale organizations, subjects had more specialized roles, e.g., developing internal security-enabling software for infrastructure. In large organizations, subjects had similar goals, but often belonged to/oversaw multiple teams.
For confidentiality, we refer to organizations by the letters A–L and participants using the notation SX(y), where $X$ denotes a unique subject number and $y$ either distinguishes participants from the same organization or represents the subject’s product type. This notation is also used to provide context to subject’s responses. For example, S10 (SAAS, Dev tools) refers to Subject S10’s direct team’s software product, which is a SaaS-type product for an organization whose product area is developer tools.
Addressing Potential Bias: While the connection to the CNCF and specific signing projects could bias responses towards certain approaches, this potential bias is balanced by the deep involvement of participants in the field of software supply chain security. Their contributions are crucial for understanding the current state and future directions of software signing practices. Moreover, the insights provided by these experts, who are at the forefront of the industry, offer a unique and valuable perspective that is essential for the study’s objectives.
Generalizability of Results: We acknowledge that the non-random sampling method and the concentration of participants from a specific initiative may limit the generalizability of our findings to the broader population of industry practitioners. However, the purpose of this study was not to achieve broad generalizability but rather to gain in-depth insights from a highly specialized group of experts—hence the use of an interpretivist analysis framework. The findings are particularly relevant to organizations and stakeholders involved in similar software supply chain security initiatives and may inform best practices and future research in this specific context. While broader generalization may be limited, the detailed, expert-driven analysis provided by this study offers significant value in advancing understanding within this niche but critical area.
@@ 4.4 Data Analysis
Interview recordings were transcribed by www.rev.com (human transcription). We anonymized identifying information. Our analysis then proceeded via multi-stage coding  [13] and a subsequent thematic  [12] and framework analysis  [72].
Three analysts (A1–A3) participated in the analysis stage of the transcripts. Analyst A1, the lead author, was responsible for all stages of data collection (including protocol design and interview conduct) and contributed to all phases of the analysis. Initial analysis was conducted by A1 and A2. When A2 became unavailable, we recruited A3. To ensure consistency, A3 was trained by memoing and coding transcripts that had already been analyzed. Comparing A3’s memos to A1 and A2, we observed strong alignment, ensuring reliability in the analysis process. A1-A3 are graduate students with research experience in cybersecurity, including software signing. The supervisors are researchers in software engineering and cybersecurity.
Multi-stage coding:
Stage 1. We began with data familiarization [77].
Stage 2. Next, we needed to develop and refine our codebook. To provide reliability while conserving resources, we used the technique described by Campbell et al. [13] and O’Connor & Joffe  [59]. First, we chose a proportion of random transcripts to multiply-code (i.e., multiple analysts coding). [^2] Although there is no common agreement on a suitable proportion of the data set to select for this, O’Connor and Joffe recommend 10–25% as typical [59]. We randomly selected approximately 33% (6 transcripts) to develop our codebook. This decision to slightly exceed O’Connor and Joffe’s recommended percentage was to familiarize analysts A2 and A3 with the nature of our collected data, proposed memo, and coding formats.
Then, two analysts (A1, A2) conducted multiple rounds of memoing on these transcripts, discussing emerging code categorizations after each round. This process resulted in 506 coded memos, which were reviewed by Analysts A1 and A3 in a series of meetings. Ultimately we obtained a codebook with 36 code categories. This codebook was further refined and extended as discussed in Stages 3-5.
Stage 3. To assess the reliability of our coding, we combine the techniques of Campbell et al. [13] and Maxam&Davis [50]. Analysts A1 and A3 randomly selected and coded a transcript independently [13, 55], and we measured their degree of agreement. Since our codebook was collectively defined by analysts A1 and A3, we supposed this coding task would not be difficult, and thus we used the percentage agreement metric as recommended by Feng [26]. In this independent coding step, we observed an 89% agreement rate. The analysts were able to resolve disagreements through discussion and code refinement. According to Campbell et al. [13], this high level of agreement suggests that a coding scheme is suitable for a single coding of the remaining uncoded transcripts.
@@
Stage 4. Next, a second set of 6 transcripts was selected and coded by Analyst A1 using the resultant codebook from Stage 3. Following Campbell et al.’s recommendation [13], we created new code categories as they emerged; this phase added 6 new code categories, each with less than 6 associated coded memos. Analyst A3 reviewed this updated codebook. Stage 5. The same process was applied to the final 6 transcripts, yielding 4 more code categories with fewer than 4 coded memos each. Of the 10 new categories recorded from stage 4 and 5, only 3 provided new insights into our data, as determined through multiple review and discussion rounds by Analysts A1 and A3. Thus A1 and A3 agreed that these new codes be classified as minor code categories.
@@
Thematic and Framework Analyses: From this process, we induced themes following Braune & Clark’s method [12]. These themes helped us answer RQs 2-4 directly. To answer RQ1 we also applied a framework analysis [72], mapping the themes to the software supply chain factory model as a reference framework to analyze participants’ responses and further categorize the themes. This framework was selected because it serves as the general reference for SSC security.
@@
Through our coding phase’s thematic analysis, we identified eight themes  Figure 6. Four of these themes were focused on the practical implementation of software signing, while the remaining four addressed various aspects of software signing: the challenges encountered, its perceived importance, the impact of security failures on its adoption, and the influence of security standards and regulations on its adoption.
@@
To determine whether additional aspects of the phenomenon might be expected to emerge if we continued to sample, we measured saturation in the 18 interview transcripts. Following Guest et al.’s [33] recommendation, we measured saturation by tracking the cumulative appearance of new codes in each interview. Each interview was rich, with a median of 33 unique codes (Figure 5). As shown in Figure 5, saturation was achieved after interview 11.
4.5 Limitation and Threats to Validity
We discuss construct, internal, and external threats to validity [92]. Following the guidance of Verdecchia et al. [84], we focus on substantive threats that might influence our findings. We also describe the role that our own perspectives and experiences may have played in this research (positionality) [24].
Construct Threats are potential limitations of how we operationalized concepts. To mitigate these threats, we grounded our interview protocol in the academic and grey literature on software signing. Following best practices, we also piloted the interview protocol internally and externally to assess its utility in capturing the desired data [15].
Internal Threats are those that affect cause-effect relationships. Qualitative data implies subjective analysis. To improve the reliability of our results, we followed an iterative process to develop the codebook and used multiple raters with a high measured level of agreement. We also acknowledge the possibility of a response bias, as some participants were recruited from the professional network of one of the authors. To address this, the involved author was not present during the interviews and played no role in the data analysis process.
External Threats may impact generalizability. This work is qualitative, which inherently carries limitations. Given the cost of an interview study, our work has a smaller sample size (N=18) than methods such as surveys can achieve. Our sample size is consistent with that of similar studies [94, 41]. Our subject recruitment approach may also have introduced biases, e.g., towards CNCF/The Linux Foundation-oriented frameworks and tools. As a trade-off, using CNCF members as a starting point improved our access to high-ranking engineering staff. To reduce bias, we also recruited through snowball sampling and other professional networks. We assessed this risk with saturation — by the 11th interview our data saturated, supporting the adequacy of our sample size (Figure 5). We also segmented our interview protocol (Table 1) and intentionally excluded the tool-specific analysis from this study. Instead, we concentrated on parts of the protocol related to the general practice of software signing.
Our study focuses on an expert population to capture nuanced insights into software signing strategies and decisions. While this approach excludes the perspectives of the broader developer community, it aligns with our objective to understand signing practices at a strategic level. Less experienced developers may interpret or prioritize signing differently, reflecting their vantage point within the software development lifecycle. By targeting experts, we ensured that the findings are informed by those directly responsible for implementing and guiding signing strategies.
@@ 4.5 Limitation and Threats to Validity
Positionality Statement: We acknowledge that our backgrounds may have influenced this study [24]. The author team are cybersecurity researchers with expertise in software signing, so our expertise both enabled and influenced the protocol development and analysis. In addition, one of the authors is a contributor to a popular signing tool. To reduce bias resulting from specific aspects of that tool, that author had no involvement in the development of the research protocol or analysis. However, that author informed the framing of the study and reviewed the results.
@@ 5.1 RQ1: Signing Implementations in Practice
Part of our interview protocol (Table 1) ascertained how subjects use software signing in their security strategies at each stage of the software supply chain factory model (Figure 1). From these responses, we identified when Software signing was required by the subject’s team (or organization) and mapped these instances back to the stages of the software supply chain factory model. The summary of our framework analysis is presented in Table 3.
Table 3: Different Signing Use Points, categorized by the stages of the software supply chain factory model highlighted in Figure 1 and subjects who employ software signing at these points. We mark subjects who consider them optional for reasons of demand by customers, nature of products, team requirements, etc. The points highlighted (PI, VI, PE, VE, PS, VS, PB, VB, PA, VA, PD, VD) are described in  Figure 3.
Points & description of Software signing use Subjects
@@
Internal code contributors sign commits-PI S2-S11, S14-S17 (S4, S8, S14–Optional)
External code contributors sign commits-PE S7, S10 (S11, S14–Optional)
Verify signatures from Internal contributors-VI S7, S14, S17
Verify signatures from external contributors-VE
Signing after code reviews/audits-PS S2-S11, S14-S17 (S4, S8, S14–Optional)
@@
Verify source code signatures before build-VS S2, S3, S7, S14
Signing of build output-PB S1-S5, S7, S8, S11-S17 (S8-Optional)
@@
Verify build signature before deployment-VB S7, S14
Signing of Final Software product-PA All S1–S18
@@
Signing after verification and certification of dependencies-PD S6, S7, S9, S16, S17, S18
Verify signatures from external dependencies-VD S9, S11, S17 (S9, S11, S17 — Optional)
Verification by final customers-VA Difficult to determine
@@ 5.1.1 Source Code
First, we ascertained whether our subjects used any form of external code contributors in their software authorship process. Out of 10 subjects who reported a form of collaboration with open-source or other external contributors, only 5 subjects used external code contributions in the production code of projects they owned (both commercial and OSS). We highlight the points and manner in which subjects reported the use of software signing to create their source code.
External contributors sign commits: Most teams who had input from external code contributors required Software signing. Out of 5 subjects (4 organizations) whose teams used external code contributions in their production code, 4 of these subjects (3 organizations) require that commits be signed. However, two of those do not require signing if the contribution is toward their open-source projects or noncritical projects. Only 2 of our subjects (2 organizations) unconditionally require signing from their external contributors. As S2 (SAAS, SSC Security tool) put it, “I got something signed from somebody I don’t know…So we do the code review, and then we have a sign-off by somebody on our team that signs off from that code”.
Internal contributors sign commits: Signing is typically required by teams for their internal code contributors. 14 of 18 subjects report that their team required commit signing from internal contributors. Three subjects (S4, S8, S14) described this as an optional requirement and largely agreed with S2’s reasoning that signing alone does not establish trust in identity. Notably, these subjects’ teams often rely on contributions from both internal and external contributors.
Verification of signatures from contributors: Although code contributions are often required to be signed, the signatures are not always verified to ensure the security and provenance of the contributions. According to our subjects, the most commonly used security methods for assessing the security of code contributions include techniques like code reviews, security audits, vulnerability scans, and tests. For example, S10 (SAAS, Dev tools) states, “I think everybody on the team does sign their commits, but we don’t really verify it…If somebody stopped signing their commits, I don’t think much would really happen.”
Code is signed after reviews/audit: The outputs of code reviews and audits are generally required to be cryptographically signed. Subjects whose teams(or organizations) mandate signing for code contributions by internal contributors also report that all activities related to code reviews and audits must be signed as well (14 subjects).